OpenStudy (anonymous):

solve from 0 to 2 pi the integral sqrt(2-2cos(t)) dt

OpenStudy (anonymous):

use the identity 2sin^2(t/2) = 1 - cos(t)

OpenStudy (anonymous):

Awesome, turns out to be pretty easy with that identity

OpenStudy (anonymous):

yep ...did you get 8?

OpenStudy (anonymous):

you should remember this identities sin^2(x) = 1/2 - 1/2 cos2x and cos^2(x) = 1/2 + 1/2 cos2x ... they are really helpful when you take integrals of even power trig functions

OpenStudy (anonymous):

(these) instead of this